$35.50 per share is the current price for Foster Farms' stock. The dividend is projected to increase at a constant rate of 5% per year. The required rate of return on the stock, rs, is 10%. What is the stock's expected price 3 years from today?
Answer #1

Today's Stock Price = \large \frac{D_{0}(1+G)}{Rate of Return- G}

35.50 = \large \frac{D_{0}(1+0.05)}{0.10- 0.05}

D0 = \large \frac{1.775}{1.05}

D0 = 1.69

Dividend at Year 0 is $ 1.69

Dividend at year end 3 or beginning of year 4 = 1.69(1+0.05)4

D4    = $ 2.054

Value of Stock 3 years from today = \bg_white \large \frac{D_{4}}{Rate of Return- Growth Rate }

= \bg_white \large \frac{2.054}{0.10- 0.05}

= 41.08

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.